# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

### v1

@@ 3 $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge(\neg\mu_{1}\mathrm{U}_{\mathrm{I}_{1}}\mu_{2})$
Table 2: STL Task Templates for Experiments
To evaluate the performance of our algorithm, we further conduct experiments in the Maze2D environment. Specifically, to validate the zero-shot generalization capability of our algorithm for STL tasks, we test it on a set of testing cases containing randomly generated STL tasks in three different Maze2D environments: U-Maze, Medium, Large. The experimental setup follows the same framework described in the Case Study. The agent starts from a randomly generated position and must complete the randomly generated STL tasks by reaching the target region within the specified time interval. All experiments were conducted on a PC running Ubuntu 22.04, equipped with an Intel i7-13700K CPU and an Nvidia 4090 GPU.
STL Tasks Generation
In order to generate random STL tasks, we design nine STL task templates as shown in Table 2. For each template, we randomly generated time intervals and the positions and sizes of circular regions corresponding to atomic predicates in the template, resulting in randomized STL tasks. Specifically, for each pair of Maze2D environment (U-Maze, Medium, Large) and each task template listed in Table 2, we generate 150 feasible random STL formulae.
Baseline Algorithm
We compare our algorithm with the method proposed in Zhong et al. (2023), which adopts classifier-based guidance and directly leverages the gradient of the trajectory’s robustness value to guide the sampling process of the diffusion model, thereby optimizing the robustness of the generated trajectory. The gradient of robustness is calculated by the STLCG method proposed in Leung et al. (2023). In the following text, we refer to this algorithm as the Robustness Guided Diffuser (RGD).
Experiment Settings
The diffusion models used in both RGD and our algorithm are trained following the procedure in Janner et al. (2022) using the D4RL dataset Fu et al. (2020). A simple multilayer perceptron (MLP) with four fully connected layers is used as the TimePredict model in our algorithm and it is also trained using the D4RL dataset. In our experiments, we employ diffusion model to generate only the state sequence of the trajectory and use a simple PD controller to follow the state sequence during running to get the actual execution trajectory.
Evaluation Metrics
For each environment and each STL task template, we test RGD and our algorithm on all randomly generated test cases and record the average of the following metrics across all cases:
@@
Env Type Success Rate(%$\uparrow$ Total Planning Time(s)$\downarrow$ T1(s)
RGD ours RGD ours
U 1 80.00 97.33 13.43$\pm$1.51 0.86$\pm$0.13 0.86
2 36.67 92.00 16.65$\pm$2.06 0.64$\pm$0.17 0.64
3 32.00 91.33 19.68$\pm$11.76 1.31$\pm$0.15 1.21
4 - 90.00 - 1.64$\pm$0.20 1.38
5 - 84.67 - 2.59$\pm$0.45 2.46
6 - 86.67 - 2.35$\pm$0.42 2.35
7 - 89.33 - 1.86$\pm$0.33 1.86
8 - 97.33 - 0.98$\pm$0.15 0.81
9 - 88.67 - 1.53$\pm$0.27 1.52
M 1 70.00 94.67 53.90$\pm$5.78 3.74$\pm$0.34 3.74
2 34.67 89.33 70.69$\pm$8.27 2.72$\pm$0.67 2.72
3 35.33 83.33 129.87$\pm$27.25 5.53$\pm$0.33 5.43
4 - 83.33 - 7.09$\pm$0.54 6.80
5 - 82.00 - 11.36$\pm$1.32 11.22
6 - 84.67 - 11.76$\pm$1.36 11.76
7 - 90.00 - 8.01$\pm$1.21 8.00
8 - 91.33 - 3.87$\pm$0.38 3.71
9 - 82.67 - 6.53$\pm$0.74 6.52
L 1 34.67 92.00 55.48$\pm$5.31 3.62$\pm$0.33 3.62
2 16.67 81.33 68.70$\pm$7.65 2.88$\pm$0.63 2.87
3 26.67 79.33 136.35$\pm$38.12 5.59$\pm$0.34 5.49
4 - 69.33 - 7.46$\pm$0.52 7.13
5 - 79.33 - 12.62$\pm$0.75 12.45
6 - 73.33 - 12.37$\pm$1.35 12.37
7 - 84.00 - 8.18$\pm$0.68 8.18
8 - 85.33 - 3.93$\pm$0.33 3.75
9 - 76.67 - 6.93$\pm$0.60 6.92
@@
Table 3: Result of Experiment in Maze2D Environment. U:U-Maze; M:Medium; L:Large; RGD: Robustness Guided Diffuser; T1:Trajectory Generation Time.
Results and Analysis
The experimental results are shown in Table 3. During testing, we found that the Robustness Guided Diffuser works only for simple STL tasks (Types 1 to 3). For more complex tasks, it faces two key issues: (1) Optimization Challenges: It is often infeasible to achieve a non-negative robustness value within a limited number of denoising steps; and (2) High Computational Cost: Complex tasks require longer trajectory generations and more complex robustness value computations, significantly increasing the computational overhead. As a result, the Robustness Guided Diffuser is suitable only for simple, short-horizon STL tasks.
In contrast, our method decomposes complex STL tasks into several short-horizon trajectory generation tasks under simpler constraints, significantly improving both planning success rate and efficiency. In relatively simple environments such as U-Maze and Medium, our algorithm achieves an actual execution success rate of over 80% across all types of STL tasks. Even in the more complex environment, Large, our algorithm maintains a success rate of at least 69% for all task templates. Additionally, the total planning time of our method is significantly lower than that of the Robustness Guided Diffuser (more than 10× faster).
By comparing the Trajectory Generation Time (T1) and Total Planning Time (T0), we identify trajectory generation as the primary efficiency bottleneck of our algorithm. For tasks with more predicates (Type 4, 9) or those involving “avoid” predicates (Type 3, 5, 6 and 7), runtime increases accordingly. This is mainly because tasks involving more predicates require longer waypoint sequences to satisfy, necessitating multiple calls to the diffusion model for trajectory generation. Additionally, for tasks involving “avoid” predicates, the use of SafeDiffuser requires solving more complex quadratic programming (QP) problems to ensure the trajectories satisfy the invariance progresses, thereby increasing the runtime. Recent advancements in accelerating the sampling process of Diffusion Models Song et al. (2020); Xiao et al. (2021); Zhou et al. (2025) could further enhance the efficiency of our approach, which will be studied in our future work.
Further Experiments
In the above experiments, since the system model in the simulation environment is unknown, we cannot rely on model-based approaches to precisely determine the feasibility of the random generated tasks. Instead, the feasibility of the STL formulas is assessed using our progress allocation module without considering the trajectory generation module. This approach may, to some extent, lead to an optimistic estimation of the success rate of our algorithm. To further evaluate the planning success rate, we utilize a custom-built environment where we have full access to both the environment information and system dynamics. In this setup, we use an optimization-based algorithm as a sound and complete solution Gilpin et al. (2020) to accurately determine the feasibility of the STL formulas. Experimental results demonstrate that the progress allocation module in our algorithm achieves a success rate of over 80% across various task scenarios, underscoring its strong completeness and robustness. Details of these additional experiments are provided in the Appendix C.
6 Conclusion
This paper introduced a hierarchical framework for trajectory planning under Signal Temporal Logic (STL) constraints. By integrating task decomposition with a diffusion model pre-trained on task-agnostic data, the proposed approach achieved high success rates, scalability, and zero-shot generalization to various STL tasks, while significantly reducing computational costs comparing to baseline algorithm. In future work, we aim to further optimize the progress allocation module by introducing uncertainty-aware time prediction and iterative optimization mechanism. These enhancements are expected to reduce the conservativeness of the algorithm and improve the solution quality. Additionally, we plan to integrate receding horizon control and accelerated diffusion sampling to further enhance the execution success rate and improve the algorithm’s efficiency.
Appendix A Related Works
A.1 STL Decomposition
To address the high complexity of STL control synthesis problems, several decomposition-based methods have been proposed Leahy et al. [2023]; Yu et al. [2023]; Kapoor et al. [2024]. In Leahy et al. [2023], the authors proposed a formula transformation-based method for multi-agent STL planning. This approach jointly decomposes an STL specification and team of agents. In Yu et al. [2023], the authors decompose STL tasks into several subtasks with non-overlapping time intervals using time interval decomposition and sequentially apply the shrinking horizon Model Predictive Control (MPC) algorithm to each short-time-interval subtask. However, this method is limited to handling STL fragments that do not include nested temporal logic operators.
The work most similar to our STL decomposition framework is Kapoor et al. [2024]. This work first decomposes STL tasks into several spatio-temporal subtasks with time variable constraints. Then, through time variable simplification, partial ordering, and slicing, the subtasks are segmented into several time intervals. Finally, a planning algorithm is subsequently used to sequentially solve the atomic tasks within each time interval. Our algorithm similarly begins by decomposing the STL task into several spatio-temporal progresses and time variable constraints. However, we adopt a search-based approach to determine the completion times and corresponding states to achieve each progress. During the search process, our method dynamically maintains the time variable constraints on-the-fly according to the completion of progresses. By incorporating the search mechanism, our algorithm achieves greater completeness compared to the incremental planning approach used in Kapoor et al. [2024]. Additionally, the dynamic maintenance of time variable constraints enables a more natural handling of relationships between subtasks and extends the applicability of our approach to more complex STL task fragments that allows “until” operator.
A.2 Planning with Diffusion Model
Recent advancements in diffusion-based planning methods highlight their remarkable flexibility, as they rely exclusively on offline trajectory datasets and do not require direct interaction with or access to the environment. By leveraging guided sampling, these methods can address a wide range of objectives without the need for retraining. This approach has been widely applied to long-horizon task planning and decision-making, facilitating the generation of states or actions for control purposes Janner et al. [2022]; Ajay et al. [2022]; Chi et al. [2023].
In the domain of diffusion-based planning for temporal logic tasks, several significant studies have been conducted Zhong et al. [2023]; Meng and Fan [2024]; Feng et al. [2024a, b]. For instance, Feng et al. [2024a] proposed a classifier-based guidance approach to direct the sampling process of the diffusion model, enabling the generation of trajectories that fulfill finite Linear Temporal Logic ($\text{LTL}_{f}$) tasks. Similarly, Feng et al. [2024b] introduced a data-driven hierarchical framework that decomposes co-safe LTL tasks into sub-tasks using hierarchical reinforcement learning. This framework integrates a diffusion model with a determinant-based sampling technique to efficiently produce diverse low-level trajectories, enhancing both planning success rates and task generalization capabilities.
For diffusion-based Signal Temporal Logic (STL) planning, Zhong et al. [2023] employed a classifier-based guidance method that leverages the gradient of robustness values to guide the sampling process of a diffusion model. This method enabled the generation of vehicle trajectories compliant with STL-specified traffic rules. Expanding on this work, Meng and Fan [2024] introduced a data augmentation process to further improve trajectory diversity and increase the satisfaction rate of specified rules. However, these approaches remain constrained to relatively simple STL tasks due to the inherent complexity of optimizing robustness values, as well as the trade-off between maximizing reward objectives and preserving the feasibility of generated trajectories Li et al. [2024].
We compare our algorithm with the method proposed in Zhong et al. [2023] in the experiments. The results demonstrate that our algorithm can handle complex and long-horizon STL tasks that the method in Zhong et al. [2023] cannot address, while also exhibiting significant advantages in runtime efficiency.
Appendix B Diffusion Models for Trajectory Planning
The core concept of diffusion-model-based trajectory planningJanner et al. [2022]; Ajay et al. [2022] is to employ the diffusion model to learn the distribution of pre-collected trajectories $q\left(\boldsymbol{\tau}^{0}\right)$ in the current environment under the system dynamics, thereby transforming planning or control synthesis problems into conditional trajectory generation.
$q(\boldsymbol{\tau}^{i}\mid\boldsymbol{\tau}^{i-1})=\mathcal{N}(\boldsymbol{\tau}^{i};\sqrt{1-\beta_{i}}\boldsymbol{\tau}^{i-1},\beta_{i}\mathbf{I}),$ (B.1)
$\boldsymbol{\tau}^{i}=\sqrt{\bar{\alpha}_{i}}\boldsymbol{\tau}^{0}+\sqrt{1-\bar{\alpha}_{i}}\boldsymbol{\varepsilon},\quad\boldsymbol{\varepsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (B.2)
with $\bar{\alpha}_{i}=\prod_{j=1}^{i}(1-\beta_{j})$.
$p_{\theta}(\boldsymbol{\tau}^{i-1}\mid\boldsymbol{\tau}^{i})=\mathcal{N}(\boldsymbol{\tau}^{i-1};\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i),\boldsymbol{\Sigma}^{i}),$ (B.3)
where $\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i)$ is parameterized by the model, and $\boldsymbol{\Sigma}^{i}$ is typically fixed. Instead of learning $\boldsymbol{\mu}_{\theta}$ directly, the model typically predicts the noise $\boldsymbol{\varepsilon}_{\theta}(\boldsymbol{\tau}^{i},i)$ and gets $\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i)$ according to the relationshipHo et al. [2020]:
$\boldsymbol{\mu}_{\theta}(\boldsymbol{\tau}^{i},i)=\frac{1}{\sqrt{\alpha_{i}}}\left(\boldsymbol{\tau}^{i}-\frac{1-\alpha_{i}}{\sqrt{1-\bar{\alpha}_{i}}}\boldsymbol{\varepsilon}_{\theta}(\boldsymbol{\tau}^{i},i)\right).$ (B.4)
The model is trained by minimizing a simplified loss function:
$\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{i,\boldsymbol{\varepsilon},\boldsymbol{\tau}^{0}}\left[\left\|\boldsymbol{\varepsilon}-\boldsymbol{\varepsilon}_{\theta}(\boldsymbol{\tau}^{i},i)\right\|_{2}^{2}\right].$ (B.5)
B.1 Safe Planning with Diffusion Model
Previous studies Xiao et al. [2023]; Botteghi et al. [2023]; Mizuta and Leung ; Zheng et al. [2024]; Christopher et al. [2024] have demonstrated the effectiveness of diffusion models in generating safety constraint-compliant trajectories. In this work, we employ SafeDiffuser Xiao et al. [2023], which integrates control barrier functions (CBFs) Nguyen and Sreenath [2016] to enforce finite-time diffusion invariance directly within the sampling process.
SafeDiffuser models the denoising process as a dynamic system:
$\dot{\boldsymbol{\tau}}^{i}=\lim_{\Delta\tau\to 0}\frac{\boldsymbol{\tau}^{i}-\boldsymbol{\tau}^{i+1}}{\Delta\boldsymbol{\tau}}=\boldsymbol{u}^{i},$ (B.6)
where $\Delta\boldsymbol{\tau}$ represents the diffusion time step, and $\boldsymbol{u}^{i}$ is a control variable with the same dimension as $\boldsymbol{\tau}^{i}$.
For each constraint $b(\mathbf{x}_{t})=h_{\mu_{i}}(\mathbf{x}_{t})\geq 0$, we define the associated CBF constraints as:
$\frac{\mathrm{d}b\left(\mathbf{x}_{t}^{i}\right)}{\mathrm{d}\mathbf{x}_{t}^{i}}\boldsymbol{u}_{t}^{i}+\boldsymbol{\alpha}(b\left(\mathbf{x}_{t}^{i}\right))\geq 0,\quad\forall i\in\{0,\dots,N-1\},$ (B.7)
where $\boldsymbol{\alpha}$ is an extended class-$\mathcal{K}$ function, and $N$ is the total number of diffusion time steps.
To ensure compliance with these constraints while minimally altering the denoising process, the optimal control $\boldsymbol{u}^{i*}$ is computed at each diffusion time step $i$ by solving the following quadratic programming (QP) problem:
$\text{minimize}\quad\left\|\boldsymbol{u}^{i}-\frac{\boldsymbol{\tau}^{i}-\boldsymbol{\tau}^{i+1}}{\Delta\boldsymbol{\tau}}\right\|^{2},$ (B.8)
subject $\text{subject to}\quad\frac{\mathrm{d}b\left(\mathbf{x}_{t}^{i}\right)}{\mathrm{d}\mathbf{x}_{t}^{i}}\boldsymbol{u}_{t}^{i}+\boldsymbol{\alpha}(b\left(\mathbf{x}_{t}^{i}\right))\geq 0.$
The updated state $\boldsymbol{\tau}^{i*}$ is then computed by substituting $\boldsymbol{u}^{i}=\boldsymbol{u}^{i*}$ into (B.6), and $\boldsymbol{\tau}^{i}$ is replaced with $\boldsymbol{\tau}^{i*}$. The complete denoising procedure is summarized in Algorithm 3.
Algorithm 3 Safe Planning with Diffusion Model
0:  Trajectory length $T$, start state $\mathbf{x}_{s}$, goal state $\mathbf{x}_{g}$, invariance constraints $\{b(\mathbf{x}_{t})\geq 0\}$
0:  A trajectory $\boldsymbol{\tau}$ of length $T$ that satisfies all constraints
1:  Initialize $\boldsymbol{\tau}^{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ with length $T$
2:  Replace $\mathbf{x}_{0},\mathbf{x}_{T}$ in $\boldsymbol{\tau}^{N}$ with $\mathbf{x}_{s},\mathbf{x}_{g}$
3:  for $i=N-1,...,0$ do
4:     $\boldsymbol{\mu}^{i+1}=\frac{1}{\sqrt{\alpha_{i+1}}}(\boldsymbol{\tau}^{i+1}-\frac{1-\alpha_{i+1}}{\sqrt{1-\bar{\alpha}_{i+1}}}\boldsymbol{\varepsilon_{\theta}}(\boldsymbol{\tau}^{i+1},i+1))$
5:     $\boldsymbol{\tau}^{i}\sim\mathcal{N}(\boldsymbol{\mu}^{i+1},\boldsymbol{\Sigma}^{i+1})$
6:     Solve the QP (B.8) and get $\boldsymbol{u}^{i*}$
7:     Calculate $\boldsymbol{\tau}^{i*}$ by setting $\boldsymbol{u}^{i}=\boldsymbol{u}^{i*}$ in (B.6)
8:     $\boldsymbol{\tau}^{i}\leftarrow\boldsymbol{\tau}^{i*}$
9:     Replace $\mathbf{x}_{0},\mathbf{x}_{T}$ in $\boldsymbol{\tau}^{i}$ with $\mathbf{x}_{s},\mathbf{x}_{g}$
10:  end for
11:  return  $\boldsymbol{\tau}^{0}$
B.2 Time-aware Trajectory Planning
As noted in Janner et al. [2022], the planning horizon of a diffusion model is not fixed by its architecture but dynamically adapts based on the input noise size. This allows for variable-length plans during the denoising process at test time. In this work, we leverage this adaptability by directly adjusting the input noise size to generate trajectories of the desired length.
During experiments, we observe that models trained with longer planning horizons generalize better to different trajectory lengths during testing. In contrast, models trained with shorter planning horizons exhibit weaker generalization. To address this, we recommend either:
  • Using variable horizons during training instead of a fixed horizon, or
  • Employing multiple models trained with different horizons to generate trajectories of diverse lengths.
Appendix C Comparative Experiment with Optimization-based Method

### v2

@@ 3 $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge(\neg\mu_{1}\mathrm{U}_{\mathrm{I}_{1}}\mu_{2})$
F.3 Details of Experiment in Maze2d Environment
Baseline Algorithm. We compare our algorithm with the method proposed in [23], which adopts classifier-based guidance and directly leverages the gradient of the trajectory’s robustness value to guide the sampling process of the diffusion model, thereby optimizing the robustness of the generated trajectory. The gradient of robustness is calculated by the STLCG method proposed in [55]. In the following text, we refer to this algorithm as the Robustness Guided Diffuser (RGD).
Experimental Settings. The diffusion models used in both RGD and our algorithm are trained following the procedure in [16] using the D4RL dataset [38]. A simple multilayer perceptron (MLP) with four fully connected layers is used as the TimePredict model in our algorithm and it is also trained using the same dataset. In our experiments, we employ diffusion model to generate only the state sequence of the trajectory and use a simple PD controller to follow the state sequence during running to get the actual execution trajectory as described in Section 3.5.
Evaluation Metrics. For each pair of Maze2D environment (U-Maze, Medium, Large) and task template listed in Table F.1, we generate 150 feasible random STL formulae as testing cases and test RGD and our algorithm on them and record the following metrics:
@@ F.3 Details of Experiment in Maze2d Environment
  • Average Robustness Value (RV): The average robustness value of the executed trajectories, after discarding the top and bottom 5% to mitigate the effect of outliers.
@@
For total planning time, we report both the mean and standard deviation in our results. For success rate, we compute the proportion of successful cases over the total number of test cases.
Full Results. The full experimental results are presented in Table F.2.
@@
Table F.2: Full Result of Experiment in Maze2D Environment. U:U-Maze; M:Medium; L:Large; RGD: Robustness Guided Diffuser; T1:Trajectory Generation Time. “-” means that RGD fails to generate feasible trajectory.
[…]
RGD ours RGD ours RGD ours
[…]
More Cases. We visualized some of the experimental results. The actual execution trajectories for some successful cases (where the STL tasks were satisfied by the execution trajectories) are shown in Figure F.2 to Figure F.4. For some failed cases (where the STL tasks were not satisfied), the trajectories planned by our algorithm are shown in Figure F.5.
Failure‑case Analysis. Notably, by analyzing the failure-cases, we identified that the primary reason for execution failure is that the trajectories generated by the diffusion model significantly violated system dynamics, such as colliding with obstacles in the environment or having excessively large distances between consecutive states. To further enhance the actual execution success rate, our method can be integrated with some receding horizon control methods [56] or online replanning strategies [57]. This extension will be explored in our future work.
Figure F.2: Successful Execution Trajectories in Maze2D-large Environment under STL Template 3: $\mathrm{F}_{\mathrm{I}_{1}}\mu_{1}\wedge(\neg\mu_{1}\mathrm{U}_{\mathrm{I}_{1}}\mu_{2})$. The task requires the agent to eventually reach region $\mu_{1}$ within a given time interval, but before reaching $\mu_{1}$, it must first visit region $\mu_{2}$.
Figure F.3: Successful Execution Trajectories in Maze2D-medium Environment under STL Template 4: $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}\wedge\mathrm{F}_{\mathrm{I}_{4}}(\mu_{4})))))$. The task requires the agent to sequentially visit regions $\mu_{1},\mu_{2},\mu_{3},\mu_{4}$ within the given time intervals.
Figure F.4: Successful Execution Trajectories in Maze2D-umaze Environment under STL Template 5: $\mathrm{F}_{\mathrm{I}_{1}}(\mu_{1}\wedge(\mathrm{F}_{\mathrm{I}_{2}}(\mu_{2}\wedge\mathrm{F}_{\mathrm{I}_{3}}(\mu_{3}))))\wedge\mathrm{G}(\neg\mu_{4})$. The task requires the agent to sequentially visit regions $\mu_{1},\mu_{2},\mu_{3}$ within the given time intervals and always avoid the region $\mu_{4}$.
Figure F.5: Planned Trajectories in Some Failure Cases.
F.4 Details of Experiments under More Complex Dynamics
To further evaluate the generality and robustness of our framework, we conduct experiments on two dynamics-rich domains from the offline goal-conditioned RL benchmark OGBench [42]: cube-single-play (“Cube”) and antmaze-medium-navigate (“AntMaze”).
Experimental Settings. The Cube environment involves a 6-DoF UR5e robotic arm manipulating a cube, while AntMaze features an 8-DoF quadruped ant navigating through a complex maze. In both environments, training relies solely on STL task-agnostic trajectory datasets provided by the benchmark; during evaluation, we replace the original goal-conditioned objectives with randomly generated STL tasks that encode multi-stage temporal and spatial constraints.
For AntMaze, we adopt the STL formulation described in Section 5.1: the agent must visit designated regions in a specific temporal order to satisfy the task specification. Planning is performed in the two-dimensional $x$-$y$ workspace, while execution uses an inverse dynamics controller that maps a 29-dimensional observation and the next $x$-$y$ target to an 8-dimensional action at each control step as in [47]. Given the higher dynamic complexity of this domain, we set $k=2$, such that each planning timestep corresponds to two control steps, as described in Section 3.5.
For the Cube environment, we focus on the end-effector’s Cartesian motion rather than the physical manipulation of objects. Trajectories are generated in $x$-$y$-$z$ space and tracked via a PD controller, as in Section 3.5. This abstraction ensures methodological consistency with other experiments while still capturing the essential planning characteristics of high-dimensional robotic control.
Evaluation Metrics. For each environment and STL task template (as defined in Table F.1), we randomly generate 100 feasible STL formulae as test cases. We report two quantitative metrics: Execution Success Rate (SR) and Total Planning Time (T0).
Results.
The complete results are summarized in Table F.3. Our framework achieves high success rates across both environments, with moderate computational cost despite the distinct dynamic structures of the two systems.
Table F.3: Results in Cube and AntMaze. SR: Execution Success Rate; T0: Total Planning Time;
Env Type SR(%) T0(s)
[…]
Analysis. In the Cube environment, our method achieves consistently strong performance, with execution success rates exceeding 90% on most task templates. The low planning time (typically below 2 seconds) indicates that the framework efficiently handles the dynamics of the 6-DoF manipulator. Performance degradation is observed only in more complex templates (Types 4 and 9), which involve long-horizon temporal dependencies and nested STL tasks. These cases require more extensive search and multiple calls to the trajectory generator, slightly increasing the planning time while reducing the success rate due to accumulated modeling uncertainty. Nevertheless, the overall high success rate demonstrates that our framework generalizes well to this manipulation-like domains.
In contrast, the AntMaze environment presents a substantially greater challenge due to its high-dimensional locomotion dynamics and strong coupling between joint configurations and global motion. Here, the average success rate remains around 80–90% for moderate templates but drops to around 60% for the most complex ones (Types 5, 6, and 9). A closer examination reveals that these harder tasks often involve stringent temporal nesting and “avoid” constraints, which are more sensitive to execution noise and controller-induced deviation. The higher planning times-especially for long-horizon templates (e.g., Type 5 and 6)-reflect the reduced efficiency of the trajectory generation module in scenarios requiring longer trajectories due to more complex dynamics. Nevertheless, our method is still able to produce feasible trajectories in most cases within an acceptable time frame.
Overall, these results confirm that the proposed framework scales effectively from smooth, fully actuated manipulation systems to complex locomotion environments. The modular combination of decomposition, allocation, and diffusion-based trajectory generation allows efficient reasoning over STL objectives, maintaining both computational efficiency and high task success rates across diverse dynamical regimes.
F.5 Details of Comparative Experiment with Optimization-based Method